A car travels 240 miles. A second car, traveling 12 mph faster that the first, makes the same trip in 1 hour less time. Find the speed of each car.
Answer by mananth(16949)

car I-----------------x mph
car II---------------x+12 mph
t=d/r
time car I -time car II =1
240/x - 240/(x+12)=1
LCD = x(x+12)
multiply by LCD
240(x+12)-240x=1*x(x+12)
240x+2880-240x=x^2+12x
x^2+12x-2880=0
x^2+60x-48x-2880=0
x(x+60)-48(x+60)=0
(x+60)(x-48)=0
x=48 mph speed of car I
48+12=60 mph Car II speed